Sam received continuous payments at an annual rate of 8t + 5 from time 0 to 10 years. The continuously compounded interest rate is 9%. Determine the present value at time 0
Answer #1

B C E 0 Year Cash flow PV factor Present value 5 1 5 1 0.9174 11.93 0.8417 17.68 0.7722 22.39 0.7084 26.21 0.6499 29.25 0.596

EXCEL FORMULA:

Year Cash flow PV factor =1/1.094B3 =C3+8 =1/1.094B4 =C4+8 =1/1.094B5 =C5+8 =1/1.09AB6 =C6+8 =1/1.09AB7 =C7+8 =1/1.094B8 =C8+